- 2 + x - 2 < 0 has no solutions. True False
Answer #1

Given statement is - x2+x-2<0 has no solution A = 32 - 4ac = (1)2 - 4 (-1)(-2) = 1-8 =-70 Hence - x2+x-2 <0 has no solution A